OpenStudy (anonymous):

Suppose that 10 grams of Carbon is released in a nuclear energy plant accident. How long will it take for the 10 grams to decay to 1 gram? (Half-Life of Carbon (14C) is 5,715 years.

OpenStudy (anonymous):

\[.1=(\frac{1}{2})^{\frac{t}{5725}}\] solve for t

OpenStudy (anonymous):

\[\frac{\ln(.1)}{\ln(.5)}=\frac{t}{5715}\] \[t=\frac{5715\ln(.1)}{\ln(.5)}\] then a calculator

OpenStudy (anonymous):

i get 18984 http://www.wolframalpha.com/input/?i=5715ln%28.1%29%2Fln%28.5%29